OpenStudy (helpmeplease14):

I don't know what to do find the nth term of each arithmetic sequence described. -9, -7, -5, -3, ...for n=18

OpenStudy (joseph127):

Just use this fomular nth = a +(n-1)*d

OpenStudy (mathmale):

"the nth, or last, term of an arithmetic sequence equals the first term plus (n-1) times the spacing between terms (d)." What is d, the spacing between terms? How big is the jump from -9 to -7?

OpenStudy (helpmeplease14):

2

OpenStudy (mathmale):

n is the index, or counter, of the terms. You could say, "seventh term" and let n = 7 represent that.

OpenStudy (mathmale):

yes, "d" is 2.

OpenStudy (mathmale):

What's the first term, a?

OpenStudy (helpmeplease14):

They all add up to 2

OpenStudy (mathmale):

The given sequence is -9, -7, -5, -3, .. What is the first term of this arithmetic sequence? That's "a". Answer this, please: a = ?

OpenStudy (helpmeplease14):

a is 2

OpenStudy (mathmale):

Actually, "d" is 2. "d" is the jump from one term of the sequence to the next. You were right on that. But look at the sequence: -9, -7, -5, -3, .. What is the first term of this sequence?

OpenStudy (helpmeplease14):

-9

OpenStudy (mathmale):

That's correct. Look at what we have now: a = -9 d = 2 n = 18 Please calculate (n-1). n-1 = ?

OpenStudy (mathmale):

Again: nth term of arith sequence = first term of sequence plus (n-1)* jump \[a _{n}=a _{1}+(n-1)d\]

OpenStudy (mathmale):

What is 18 - 1 -------- ??

OpenStudy (helpmeplease14):

17

OpenStudy (helpmeplease14):

Don't you do distributive property???

OpenStudy (mathmale):

You could, but it's easier simply to recognize that 18-1=17 and then multiply this (n-1) by d, or to multiply this 17 by 2. What is 17 * 2?

OpenStudy (helpmeplease14):

34

OpenStudy (mathmale):

And the first term is -9. The 18th term is then -9 + 34. Please calculate this.

OpenStudy (mathmale):

\[a _{18}=-9+34=?\]

OpenStudy (helpmeplease14):

Why is it a18 and not -9???

OpenStudy (helpmeplease14):

Got 25 btw

OpenStudy (mathmale):